Mr. Domingo drives to Laguna for the weekend at an average speed of 105km/h and the trip takes 2hrs and 20mins. On the way back to Manila, he took the same road (covering the exact same distance) but heavy traffic slows him down to an average speed of only 70km/h.

a) How many hours did it take him on the way back?


b) By how much longer is the trip back compared to the first trip?




1
Expert's answer
2021-03-29T09:03:53-0400

"t_1=2^h20^m=\frac73 ~hours,"

"s=v_1t_1=v_2t_2,"

a)

"t_2=t_1\cdot \frac{v_1}{v_2}=\frac73\cdot \frac{105}{70}=3.5~hours,"

b)

"\Delta t=t_2-t_1=3.5-\frac 73=1^h10^m ."